OpenStudy (anonymous):

PLEASE: Determine if the graph of f(x)=x^2+x if ≠0. and 0 if x=0 is continuous for all real numbers, state “no discontinuities” and explain why using the conditions of continuity. If there are any discontinuities, please: a-Determine what x value the discontinuity occurs b-Determine if the discontinuity if removable or non-removable c-If removable, identify the value of the hole. If non-removable, show proof using the conditions of continuity.

OpenStudy (anonymous):

I just need a, b and c PLEASEEE

OpenStudy (zepp):

\(y=x^2\) isn't continuous when \(x=0\)

OpenStudy (anonymous):

so that is a? @zepp

OpenStudy (anonymous):

can you do b and c then?

OpenStudy (zepp):

Oh, \(x^2+x\), my bad, have to check again D:

OpenStudy (anonymous):

ok

OpenStudy (zepp):

Ok, a removable discontinuity occurs when the limit from left and right are the same, but from one side, it's undefined.

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I have been trying forever, can u just please tell me a,b,c? :)

OpenStudy (zepp):

:D

OpenStudy (anonymous):

So what would I put??

OpenStudy (anonymous):

I understand that, but what I would i put for a b and c @zepp and @apoorvk

OpenStudy (zepp):

Simply verify the limit from 0+ and 0-, and well, when it's undefined, it's at this point the discontinuity occurs.

OpenStudy (apoorvk):

c'mon, it's on the tray now. I toldya the function is non-continuous at (0,0). So part (a) is staring at your face. But please try to understand the procedure, or you won't be able to solve the other parts. (and am sorry but simply 'answering' is not what we help with, though sure as hell we can toil along with you :))

OpenStudy (anonymous):

Ok let me see if I can answer then and tell me if I'm right?

OpenStudy (anonymous):

a) x=0 ?

OpenStudy (anonymous):

b) removeable and c) 0

OpenStudy (anonymous):

@apoorvk @zepp is that right???

OpenStudy (apoorvk):

You mean (0,0) right for C? yeah they seem okay to me! Btw just confirming in the first line of the question, what's " ≠0. " ??

OpenStudy (anonymous):

does not equal 0

OpenStudy (zepp):

I don't think it's removable D:

OpenStudy (anonymous):

is it removable or not? I'm not sure

OpenStudy (zepp):

There's no other point to repair that hole right there.

OpenStudy (anonymous):

@apoorvk what do u think? removable or no

OpenStudy (anonymous):

this was my answer: Discontinuous at (0,0)! a) x=0 b) removable c) (0,0) is the hole.

OpenStudy (anonymous):

is that right

OpenStudy (apoorvk):

OMG. Does the question mean the function is not equal to zero, but is 0 for x=0 only. Does it? @zepp and @schmidtdancer

OpenStudy (anonymous):

yes it says ≠..

OpenStudy (zepp):

@_@ I'm confused D:

OpenStudy (anonymous):

so is my answer right??

OpenStudy (anonymous):

it says ≠

OpenStudy (anonymous):

@apoorvk @zepp

OpenStudy (apoorvk):

I think, thanks to the very stupid language of the question combined with my own stupidity, we may have to solve this again totally. Sorry folks :(

OpenStudy (anonymous):

really? so my answer isn't right?!!?

OpenStudy (zepp):

I don't understand this part f(x)=x^2+x if ≠0. and 0 if x=0 If x≠0?...and x=0.. what?

OpenStudy (anonymous):

I mean f(x)=x^2+x if x≠0. and 0 if x=0

OpenStudy (anonymous):

so @zepp @apoorvk Idk....is my answer ok?

OpenStudy (apoorvk):

Oh I typed everything and misunderstood it again. Anyways, i'm got him now! :D So, yeah, when x=0, f(x)=0 as well. right?

OpenStudy (anonymous):

Yes! so that makes my answer right, correct????

OpenStudy (apoorvk):

And, \(x^2+x\) is itself '0' at x=0, right?

OpenStudy (apoorvk):

Forget the answer, just listen to me for a moment -.-

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

So, the function graph is actually: